Science Report – May 1st

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters.

2. Sunspot and Solar Flare Monitoring: One small, shallow prominence was photographed on the Sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water. Using dirt to clean the dishes seems to have some rapid growing bacteria!

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported. One helmet fogged today on EVA during a brief rainfall. Once rain stopped, it returned to un-fogged almost immediately.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Science Report – April 30th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters. The second EVA found a location in Candor Chasma that had been previously scouted by the other EVA team last week.

2. Sunspot and Solar Flare Monitoring: The crew astronomer saw one filament on the sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Science Report – April 29th

Science Report for 29 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no sites were noted as useful for emergency evacuation shelters. If a crew was trapped north on Copernicus highway, there are several low sites where they could shelter out of the wind.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found two dark features on the sun and another possible solar flare.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: The first EVA of the day took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Final samples of the project were taken from a towel wiped dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria growth and monitoring will continue over the remaining days of the mission.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No fogging was reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Thanks,
Doug
Crew 193 XO and Science Officer

Science Report – April 28th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs took place today.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar prominence during his imaging session this morning.

3. Shortwave Texting on EVA’s: No EVAs took place today.

4. GPS Route Measurements: No EVAs took place today.

5. Waterless dish cleaning: Samples were taken from a dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria are growing on the samples from the first test.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No EVAs took place today.

8. Crew Comfort, Health and Safety on EVA: No EVAs took place today.

Science Report – April 27th

Science Report for 27 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No locations were scouted during the EVA today. Emergency triage procedures were practiced so that a crew member might be safely transported to an evacuation location in the event of an incident.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar flare in the direction of our planet and immediately alerted mission control to the danger.

3. Shortwave Texting on EVA’s: Hand signal communication was used during the emergency EVA practice today. Further practice is needed to achieve a smooth process.

4. GPS Route Measurements: No measurements today, as the EVA was at the hab location.

5. Waterless dish cleaning: Samples were taken from a dirty dish and a dirty dish that had been put under the UV light. There are minimal signs of bacterial growth from the first set of samples. It may take a couple more days to see signs of growth. Both sets of samples are now in the incubator in the hab.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Analysis will be done as the mission continues.

Science Report – 26 April 2018

1. Emergency EVAC EVA: No locations were scouted during either EVA today. Planning for an EVA dedicated to emergency procedures is in the works.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. The sunspots are getting smaller – they were not visible with the eye but could still be seen with the Skyris camera.

3. Shortwave Texting on EVA’s: Beartooth devices still not working. Hand signal communication was used instead.

4. GPS Route Measurements: GPS was taken on EVA 8 this morning which proved very helpful in re-locating the rover after exiting the moon valley area. A picture was taken of the GPS device when EVA 8 parked for exploration and the crew the followed the GPS back to that set of coordinates.

5. Waterless dish cleaning: No samples were taken today. No growth seen on samples from previous day. Previous day samples were put into the incubator in the science hab today to speed bacterial growth.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today. One crew member thought their mask had fogged, but it turned out to just be dirt.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working very hard (normally climbing a large hill). Gold 01 is able to track his heart rate over the entire course of the EVA thanks to mobile technology. Results can be exported and used in this study. Analysis will be done as the mission continues.

Science Report April 25th

Hello Mission Support,

Here is the Science Report for 25 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by CGA Eric Shear

1. Emergency EVAC EVA: EVA 6 did not find any suitable locations for evacuation in Lith Canyon. The distance is very far from the hab, difficult to reach and quite exposed to the environment. Although it is beautiful! EVA 7 was to walk south alongside the Hab Ridge Road on foot, but was cut short due to warm conditions. Further investigation of possible locations will be done on future EVAs.

2. Sunspot and Solar Flare Monitoring: The sunspot from the previous observation has split into three smaller ones, and one prominence was seen. For photos, please refer to the astronomy report.

3. Shortwave Texting on EVA’s: After troubleshooting, one Beartooth device continues to not work today. We are awaiting tech support from the manufacturer. In the meantime, we continue to use the note app on our phones to communicate with our deaf crew member.

4. GPS Route Measurements: A GPS tracking route was successfully taken on EVA 7 with the Garmin GPSmap 60Cx. On this EVA, we walked 641 meters (roundtrip).

5. Waterless dish cleaning: Baseline testing was completed this afternoon. 2 swabs were taken from dirty dishes in the sink and two swabs were taken from dishes in the cupboard. The swabs were rubbed on a water/agor mix in petri dishes and will be watched over the next few days for bacteria growth.

Science Report – April 24th

Here is the Science Report for 24 April 2018

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Both EVAs today had objectives of identifying more sites which would be possible shelters during an emergency evacuation. EVA 4 was able to identify a cave on the south side of west kissing camels ridge which would be an excellent shelter for all possible issues. EVA 5 located the shelter site that was proposed on EVA 1 on the north ridge. EVA 5 also located another possible evacuation site in the Tank Wash area. Further investigation of possible locations will be done on future EVAs.

2. Sunspot and Solar Flare Monitoring: Please refer to the astronomy report.

3. Shortwave Texting on EVA’s: Beartooth devices did not work today. Will need to troubleshoot the devices this evening. Hand sign language was used instead to communicate with Eric as backup means of communication was left on an ATV.

4. GPS Route Measurements: GPS was taken on EVA 5 which has just returned to the hab. Data will be analyzed for a more fulsome report tomorrow.

5. Waterless dish cleaning: No progress today. Initial testing of dirty dish bacteria growth will commence tomorrow with a plan to do testing on every second day due to the amount of supplies brought by gold XO for the research.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working very hard (normally climbing a large hill). Gold 01 is able to track his heart rate over the entire course of the EVA thanks to mobile technology. Results can be exported and used in this study. Analysis will be done as the mission continues.

Science Report – April 20th

MDRS Crew 192 Final Science Report (We science the sh*t out of this!)
Richard Blakeman, Executive Officer
April 20, 2018 (Sol 12)

MDRS Crew 192 Mission Science report summary:
This report describes the outcomes of the science and research projects conducted during the MDRS Blue Crew 192 mission. It should be noted that on some of the research studies, the primary data collected requires additional time for analysis and conclusions.

Spacesuit visor fogging study
This research was conducted using a double-blind study to test off the shelf cleaning products (Joy dishwashing liquid and Johnson’s Baby shampoo) and their effectiveness against spacesuit visor fogging. Both one-piece and two-piece (separate helmet) spacesuit configurations were tested along with random controls to identify variables and collect data. The data will require analysis before any final conclusions can be made; however, preliminary data suggests that exertion levels contribute to fogging phenomenon. Additionally, baby shampoo appears to have slightly superior results in fog reduction. There were occasional reports of minor irritation but it appears to be not significant.

Hand exercises using hand relief, well-being balls
This research was conducted as single blind study to test the use of well-being ball for had exercise before EVA determining the dexterity and comfort of hands. After few measurements, discontinued the study as the exercises were creating discomfort for the crew

Crew wellness observations This is survey-based study using the Wellbeing questionnaire before, during and the end of the study to measure the happiness scale of the crew

Crew weight measurements and analysis (EVA) Daily weight measurements were taken along with the pre and post EVA analysis. Preliminary results indicated the weight loss after EVA is proportional with duration of EVA and physical exertion

Crew muscle measurements Daily crew skeletal measurements including deltoid and calf muscles were taken. Preliminary analysis show reduction in deltoid muscle in majority of the crew.

MAG (Maximum Absorbency Garment) utilization (MAG, i.e. Depend-type undergarments) were worn by crewmembers on all EVAs. This provided additional crew comfort, health, welfare, and safety protection on increasingly longer and complex EVAs. NASA and the military use the MAG protocol for missions involving extended operations involving pressure suits, EVA space suit, and undersea hard suits where waste collection issues can significantly impact crew heath and mission success.

Ultrasonic rodent repulsion experiment Three off the shelf plug-in ultrasonic rodent repulsion emitters were placed in the lower habitat, crew quarters deck, and the upper level deck. There was only one intrusion of a rodent during the mission located on the crew living deck near the refrigerator. A trap was baited with a small piece of bread coated with peanut butter and the intruding rodent was captured unharmed. On a subsequent EVA the rodent was released on Galileo Road (Route 1104). An additional rodent intruder was discovered during the night in the south-side, upper level, interface between the wall and the habitat roof structure. The intruder rodent was caught in a glue trap and did not survive. The initial conclusion is that the ultrasonic rodent repulsion emitters are ineffective. Physical traps need to be continuously deployed to capture intruder rodents and additional repulsion technologies tested.

Astronomy discussions and visual observations Conducted night time observational astronomy lectures describing various constellations and planets. The crew was able to observe several satellites and wonder at the incredible view of the heavens above. Additionally, conducted daytime solar observations using the MDRS solar telescope array. However, computer interface issues and some clouds affected viewing. Some imagery was obtained using the optical sun lens and a smart phone.

Geology observations conducted during EVAs Each EVA offered a rich and immersive experience into the local geology. Close physical inspection of structures as well as photographic and video imagery was taken for later discussion and analysis.

EVA touch screen glove testing The crew brought several types of touch pad sensitive gloves to use during EVAs. These proved to be an invaluable tool for the crews as it allowed direct interface with multiple electronic recording devices. Recommend that these be used by future crew to assist with video and photographic imagery.

Water contamination prevention and mitigation procedures
All of the habitat water storage tanks were meticulously cleaned and sanitized over the course of many days to remove any contamination and tank residue; additionally, multiple fresh water transport and loading runs to and from Hanksville was accomplished by the crew. The water transfer pump was also meticulously clean to prevent future contamination. The main water filter was also replaced by the crew.

Yuri’s night distilled spirits experiment The distillation of a celebratory spirit was both a crew morale booster and a fascinating chemistry experiment. The process took several days to complete and the resulting product was equally distributed to each crewmember in a celebratory toast to the accomplishments of Cosmonaut Yuri Gagarin for becoming the first human in space April 12, 1961.

Still and video imagery Still and video imagery has been collected by all crewmembers throughout the mission in order learn and better appreciate the challenges and requirements necessary to be an effective Martian crew.

Spacesuit hydration prototype system operational testing and evaluation An experimental prototype EVA hydration system was constructed and operationally tested on multiple EVAs both mounted and dismounted. This system has shown promising results as it can be utilized while operating a rover, ATV, as well as dismounted EVAs. Astronaut hydration, particularly during heavy exertion, is an important physiological need and critical to crew safety.

Warm Regards from Florida,

Joseph Dituri, PhD (c), CDR, US Navy Saturation Diving Officer (ret)